# On Discovering Algorithms for Adversarial Imitation Learning

### v1

@@ Abstract
Adversarial Imitation Learning (AIL) methods, while effective in settings with limited expert demonstrations, are often considered unstable. These approaches typically decompose into two components: Density Ratio (DR) estimation $\frac{\rho_{E}}{\rho_{\pi}}$, where a discriminator estimates the relative occupancy of state-action pairs under the policy versus the expert; and Reward Assignment (RA), where this ratio is transformed into a reward signal used to train the policy. While significant research has focused on improving density estimation, the role of reward assignment in influencing training dynamics and final policy performance has been largely overlooked. RA functions in AIL are typically derived from divergence minimization objectives, relying heavily on human design and ingenuity. In this work, we take a different approach: we investigate the discovery of data-driven RA functions, i.e, based directly on the performance of the resulting imitation policy. To this end, we leverage an LLM-guided evolutionary framework that efficiently explores the space of RA functions, yielding Discovered Adversarial Imitation Learning (DAIL), the first meta-learnt AIL algorithm. Remarkably, DAIL generalises across unseen environments and policy optimization algorithms, outperforming the current state-of-the-art of human-designed baselines. Finally, we analyse why DAIL leads to more stable training, offering novel insights into the role of RA functions in the stability of AIL. Code is publicly available: https://github.com/shshnkreddy/DAIL.
footnotetext: Correspondence to shshnkreddy@gmail.com.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved impressive results across a range of complex domains(Mnih et al., 2015; Silver et al., 2016), conditioned on the availability of well-defined and informative reward functions. However, in many real-world settings, specifying such reward functions is either prohibitively difficult or entirely infeasible, whereas providing demonstrations of the desired behavior is often easier and cost-effective. This motivates the paradigm of Imitation Learning (IL) (Argall et al., 2009; Schaal, 1999), which seeks to learn policies directly from expert demonstrations. IL is particularly well-suited to applications such as autonomous driving (Pomerleau, 1988) and robotic manipulation (Argall et al., 2009), where hand-crafting precise reward functions poses a significant challenge.
@@
A notably effective approach within imitation learning is Adversarial Imitation Learning (AIL) (Ho & Ermon, 2016), which draws inspiration from Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and is recognized for its strong performance when expert demonstrations are limited. Similar to GANs, AIL formulates the learning process as a two-player adversarial game between a generator (i.e., the policy network) and a discriminator network. The policy aims to generate trajectories that are indistinguishable from those of the expert, while the discriminator learns to distinguish between expert and policy-generated trajectories.
@@
Since it shares a similar objective to GANs, AIL inherits some of the training challenges associated with adversarial methods (Arjovsky et al., 2017), most notably issues related to instability. In this work, we focus on one critical factor underpinning stable and effective training: the quality of the learning signal. Prior research in GANs (Goodfellow et al., 2014; Arjovsky et al., 2017) has shown that providing strong and informative gradient signals to the generator is essential for improving its performance and ensuring convergence of the adversarial game. In the context of AIL, this learning signal manifests as the rewards given to states-action pairs visited by the policy. While recent non-adversarial approaches (e.g., (Kostrikov et al., 2020; Garg et al., 2021)) have shown promise, their empirical performance remains mixed (Jain et al., 2024; Lai et al., 2024), and offer less flexibility (eg. for reward shaping (Sapora et al., 2024)), motivating the need to stabilize adversarial imitation learning through more informative reward signals for policy optimization.
@@
We first introduce AIL and reward assignment through the lens of divergence minimization between expert and policy occupancy measures as highlighted in (Ghasemipour et al., 2020). This perspective reveals a natural two-stage decomposition of the reward assignment process: (a) Density Ratio (DR) estimation, where the ratio of occupancy measures for each state-action pair is estimated, and (b) Reward Assignment (RA), which maps this ratio to scalar rewards for policy optimization. While prior work has improved stage (a) through better stabilization of the discriminator training (Luo et al., 2024; Wang et al., 2024; Lai et al., 2024) (b) has attracted considerably less attention (Fu et al., 2018; Ghasemipour et al., 2020). Following Ghasemipour et al. (2020), we highlight how RA functions (Figure 1) influence policy learning dynamics in adversarial training, and propose an LLM-based meta learning framework to discover reward assignment functions for improved performance.
@@
This optimization produces a RA function which, when integrated into the AIL framework, results in Discovered Adversarial Imitation Learning (DAIL) (Figure 1). When evaluated on unseen environments from the Brax (Freeman et al., 2021) and Minatar (Young & Tian, 2019) suites, DAIL outperforms state-of-the-art baselines, including GAIL (Ho & Ermon, 2016), AIRL (Fu et al., 2018), FAIRL (Ghasemipour et al., 2020) and GAIL-heuristic (Orsini et al., 2021). To the best of our knowledge, DAIL is the first meta-learned AIL algorithm. We further demonstrate that it also generalizes to policy optimization algorithms not seen during discovery. Finally, by examining DAIL’s training dynamics, we show that DAIL enhances performance by producing more informative learning signals.
@@ 2 Related Work
In settings where reward design is challenging (Argall et al., 2009) or exploration is hard (Nair et al., 2018), learning from expert demonstrations offers a compelling alternative. The simplest approach, Behavior Cloning (BC) (Pomerleau, 1988; Torabi et al., 2018), treats imitation as supervised learning but struggles in low-data regimes due to compounding errors outside the training distribution. In contrast, distribution-matching methods explicitly align the expert and policy state-action distributions, which helps mitigate distributional shift and improves robustness. Among them, Adversarial Imitation Learning (Ho & Ermon, 2016; Fu et al., 2018; Ghasemipour et al., 2020; Orsini et al., 2021)—inspired by GANs—has shown strong results but suffers from instability. Prior work has focused on stabilizing the discriminator via improved loss functions (Luo et al., 2024; Wang et al., 2024), architectures (Lai et al., 2024), and regularizers (Orsini et al., 2021; Xiao et al., 2019). However, reward assignment—the mapping from discriminator logits to the reward—has received little attention. Fu et al. (2018) first highlighted the impact of RA, showing that performance can vary with the structure of the underlying MDP, and proposed methods such as incorporating absorbing states to mitigate this dependence. Ghasemipour et al. (2020) highlight the impact of RA functions on policy optimization and unify various RA functions as instances of divergence minimization within the distribution-matching IL framework, while Orsini et al. (2021) empirically benchmark these functions across multiple tasks. This work is, to the best of our knowledge, the first that explores the discovery of novel reward assignment functions, as an alternative to relying on the existing human-derived ones. Recent non-adversarial distribution-matching IL methods, such as ValueDICE (Kostrikov et al., 2020) and IQ-Learn (Garg et al., 2021), have demonstrated promise but show mixed empirical performance, with AIL variants performing comparably or better across multiple benchmarks (Lai et al., 2024; Jain et al., 2024). Additionally, unlike AIL approaches, they provide limited flexibility in accommodating scenarios such as state-only demonstrations (Torabi et al., 2018; Jain et al., 2024) and reward shaping (Sapora et al., 2024). Another closely related area is Inverse Reinforcement Learning (Ng & Russell, 2000), which seeks to infer the underlying reward function that best explains the expert policy (Skalse & Abate, 2023; Chirra et al., 2024). This contrasts with our narrower focus on directly optimizing imitation policies and designing reward assignment functions that ensure their stable training. On the other hand, works such as (Ma et al., 2024a; b; Li et al., 2025) directly tackle the problem of reward design in MDPs by leveraging LLM-based evolution, ultimately outperforming human-designed rewards.
@@
Meta-Learning (also known as ‘learning to learn’) aims to automatically discover learning algorithms through end-to-end optimization (Schmidhuber, 1987; Thrun & Pratt, 1998) and has seen extensive application in reinforcement learning (RL) (Beck et al., 2025). Historically, RL training pipelines have been limited by CPU-bound environments, however, the advent of GPU-accelerated environments has enabled up to $1000\times$ speedups—revitalizing Meta-RL. This has subsequently facilitated the discovery of novel loss functions (Lu et al., 2022; Jackson et al., 2024), activation functions (Nadimpalli et al., 2025), and optimizers (Goldie et al., 2024) for Deep RL. Differentiating through the learning process of an RL algorithm remains a challenging problem, which has motivated the use of black-box optimization methods (Salimans et al., 2017; Chen et al., 2023b). However, these approaches typically suffer from high sample complexity and limited interpretability. Large language models (LLMs) provide a promising alternative, leveraging broad domain knowledge and code-generation capabilities to produce more interpretable and effective learning algorithms (Goldie et al., 2025). Beyond RL, LLM-based black box optimization has demonstrated success across a wide range of domains, including environment generation (Faldor et al., 2024), neural architecture search (Chen et al., 2023a), combinatorial optimization (Ye et al., 2024), mathematics (Romera-Paredes et al., 2024), and more (Novikov et al., 2025).
@@ 3 Background
We consider a Markov Decision Process (MDP) $\mathcal{M}$ defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,\mu)$, where $\mathcal{S}$ denotes the set of states, $\mathcal{A}$ is the set of actions, $\mathcal{P}(s^{\prime}|s,a)\in[0,1]$ is the transition probability, $r(s,a)\in\mathbb{R}$ is the reward function, $\gamma\in[0,1]$ is the discount factor and $\mu(s)\in\Delta(\mathcal{S})$ is the initial state distribution. A policy $\pi(.|s)\in\Delta(\mathcal{A})$ is a distribution over the set of valid actions for state $s$. A trajectory $\tau=\{(s_{t},a_{t})\}$ denotes the state-action pairs encountered by executing $\pi$ in $\mathcal{M}$. For a given policy $\pi$, the occupancy measure $\rho_{\pi}(s,a)$ is defined as $\rho^{\pi}(s,a)=(1-\gamma)\mathbb{E}_{\tau\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s,a_{t}=a)$]. Intuitively, it can be interpreted as the distribution over state-action pairs that the agent encounters while following policy $\pi$. A one-to-one correspondence exists between $\pi$ and $\rho_{\pi}$ (Syed et al., 2008), allowing us to use them interchangeably. For functions $f(s,a)$ dependent only on state-action pairs, we have $\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}f(s_{t},a_{t})\right]=\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim\rho_{\pi}}[f(s,a)].$ We leverage this identity to use the two expectations interchangeably when optimizing over $f$.
@@
Jensen-Shannon GAIL (Ho & Ermon, 2016) $\text{softplus}(\ell(s,a))$
@@ 4 Problem Definition
In this paper, we focus on Step 3, highlighting that providing an informative learning signal is crucial for effective policy improvement and overall adversarial training. Figure 1 illustrates how the reward assignment function shapes the policy’s learning dynamics. The AIRL RA function encourages the policy to visit state-action pairs where expert visitation exceeds its own and penalizes it equally when it surpasses the expert. In contrast, the GAIL RA function only incentivizes matching the expert on underrepresented pairs, while its heuristic variant does the opposite. FAIRL employs a more nuanced approach: it rewards the policy for slightly exceeding expert visitation but imposes steep penalties when the expert dominates.  (Ghasemipour et al., 2020) reason this drives the policy to gradually expand and cover the expert distribution from the outside in. Furthermore, although existing RA functions (Table 1) derive from well-established $f$-divergence theory, they neglect the practical stability challenges that arise during training.
This raises the question: Can we meta-learn a reward assignment function that results in stable and effective adversarial training? To evaluate training quality, we measure the divergence between the expert and the policy after training. Specifically, we use the Wasserstein distance (Rubner et al., 1998), computed between rollouts generated by the policy and expert demonstrations. We select this metric due to its robustness and sensitivity in measuring distances between occupancy distributions in RL settings (Luo et al., 2023; Rupf et al., 2024).
@@ 5 Discovering RA functions via Evolutionary Search
Optimizing Equation 3 is challenging because it requires backpropagating gradients through the entire adversarial training loop, which is generally computationally intractable. Consequently, prior work has typically relied on black-box methods for such bilevel optimization problems (Goldie et al., 2024; Lu et al., 2022). In this work, we adopt an LLM-guided evolutionary framework—shown to be sample-efficient, interpretable, and capable of discovering generalizable algorithms in meta-RL (Goldie et al., 2025).
@@ 6 Empirical Studies
We begin by outlining the experimental setup, after which we present the results. We conduct our experiments on two benchmark suites: MuJoCo control tasks (Todorov et al., 2012) (Ant, Reacher, Walker2d, HalfCheetah, and Hopper) and Minatar (Young & Tian, 2019) (Asterix, SpaceInvaders, and Breakout). For each task, we collect 10 successful expert demonstrations from a PPO-trained policy (Schulman et al., 2017), and subsample every 20th transition—following standard practice (Ho & Ermon, 2016). Unless stated otherwise, we optimize the policy using PPO and regularize the discriminator with a gradient penalty (Gulrajani et al., 2017). Consistent with prior work in IL, we adopt fixed-length episodes of 1000 timesteps for MuJoCo environments (Gleave et al., 2022). However, we do not enforce this for the Minatar tasks since doing so significantly degraded performance. Our implementation is fully written in JAX (Bradbury et al., 2018), using PureJaxRL (Lu et al., 2022), Brax (Freeman et al., 2021) for MuJoCo environments, Gymnax (Lange, 2022) for Minatar, and OTT-JAX (Cuturi et al., 2022) for computing Wasserstein distance. Full training hyperparameters are provided in Appendix D. We normalize returns using min-max scaling between random and expert policy performance, with all results averaged over 16 independent seeds.

### v2

@@ Abstract
Adversarial Imitation Learning (AIL) methods, while effective in settings with limited expert demonstrations, are often considered unstable. These approaches typically decompose into two components: Density Ratio (DR) estimation $\frac{\rho_{E}}{\rho_{\pi}}$, where a discriminator estimates the relative occupancy of state-action pairs under the policy versus the expert; and Reward Assignment (RA), where this ratio is transformed into a reward signal used to train the policy. While significant research has focused on improving density estimation, the role of reward assignment in influencing training dynamics and final policy performance has been largely overlooked. RA functions in AIL are typically derived from divergence minimization objectives, relying heavily on human design and ingenuity. In this work, we take a different approach: we investigate the discovery of data-driven RA functions, i.e, based directly on the performance of the resulting imitation policy. To this end, we leverage an LLM-guided evolutionary framework that efficiently explores the space of RA functions, yielding Discovered Adversarial Imitation Learning (DAIL), the first meta-learnt AIL algorithm. Remarkably, DAIL generalises across unseen environments and policy optimization algorithms, outperforming the current state-of-the-art of human-designed baselines. Finally, we analyse why DAIL leads to more stable training, offering novel insights into the role of RA functions in the stability of AIL.
footnotetext: Corresponding authorfootnotetext: Research conducted while at SMU as a Research Engineer
@@ 1 Introduction
Reinforcement Learning (RL) has achieved impressive results across a range of complex domains (Mnih et al., 2015; Silver et al., 2016), conditioned on the availability of well-defined and informative reward functions. However, in many real-world settings, specifying such reward functions is either prohibitively difficult or entirely infeasible, whereas providing demonstrations of the desired behavior is often easier and cost-effective. This motivates the paradigm of Imitation Learning (IL; Argall et al. (2009); Schaal (1999)), which seeks to learn policies directly from expert demonstrations. IL is particularly well-suited for applications such as autonomous driving (Pomerleau, 1988) and robotic manipulation (Argall et al., 2009), where hand-crafting precise reward functions poses a significant challenge.
@@
A particularly effective approach within imitation learning is Adversarial Imitation Learning (AIL; Ho and Ermon (2016)), which draws inspiration from Generative Adversarial Networks (GANs; Goodfellow et al. (2014)) and is recognized for its strong performance when expert demonstrations are limited. Similar to GANs, AIL formulates the learning process as a two-player adversarial game between a generator (i.e., the policy network) and a discriminator network. The policy aims to generate trajectories that are indistinguishable from those of the expert, while the discriminator learns to distinguish between expert- and policy-generated trajectories.
@@
Since it shares a similar objective to GANs, AIL inherits some of the training challenges associated with adversarial methods (Arjovsky et al., 2017)—most notably the issues related to instability. In this work, we focus on one critical factor that underpins stable and effective training: the quality of the learning signal. Previous research in GANs (Goodfellow et al., 2014; Arjovsky et al., 2017) has shown that providing strong and informative gradient signals to the generator is essential for improving its performance and ensuring convergence of the adversarial game. In the context of AIL, this learning signal manifests as the rewards given to states-action pairs visited by the policy. Although recent non-adversarial approaches (e.g., Kostrikov et al. (2020) and Garg et al. (2021)) have shown promise, their empirical performance remains mixed (Jain et al., 2024; Lai et al., 2024), and they offer limited flexibility (e.g. for reward shaping (Sapora et al., 2024)), motivating the need to stabilize adversarial imitation learning through more informative reward signals for policy optimization.
@@
We first introduce AIL and reward assignment through the lens of divergence minimization between expert and policy occupancy measures as highlighted in Ghasemipour et al. (2020). This perspective reveals a natural two-stage decomposition of the reward assignment process: (a) Density Ratio (DR) estimation, where the ratio of occupancy measures for each state-action pair is estimated, and (b) Reward Assignment (RA), which maps this ratio to scalar rewards for policy optimization. While prior work has improved stage (a) through better stabilization of the discriminator training (Luo et al., 2024; Wang et al., 2024; Lai et al., 2024) (b) has attracted considerably less attention (Fu et al., 2018; Ghasemipour et al., 2020). Following Ghasemipour et al. (2020), we highlight how RA functions (Figure 1) influence policy-learning dynamics in adversarial training, and propose an LLM-based meta learning framework to discover reward assignment functions for improved performance.
@@
This optimization produces a RA function that, when integrated into the AIL framework, results in Discovered Adversarial Imitation Learning (DAIL) (Figure 1). When evaluated on unseen environments from the Brax (Freeman et al., 2021) and Minatar (Young and Tian, 2019) suites, DAIL outperforms state-of-the-art baselines, including GAIL (Ho and Ermon, 2016), AIRL (Fu et al., 2018), FAIRL (Ghasemipour et al., 2020) and GAIL-heuristic (Orsini et al., 2021). To the best of our knowledge, DAIL is the first meta-learned AIL algorithm. We further demonstrate that it also generalizes to policy optimization algorithms not seen during discovery. Finally, by examining DAIL’s training dynamics, we show that DAIL enhances performance by producing more informative learning signals.
@@ 2 Related Work
In settings where reward design is challenging or limited in expressivity (Argall et al., 2009; Teoh et al., 2025), or where exploration is difficult (Nair et al., 2018), learning directly from expert demonstrations offers a compelling alternative. The simplest approach, Behavior Cloning (BC; Pomerleau (1988); Torabi et al. (2018)), treats imitation as supervised learning but struggles in low-data regimes due to compounding errors outside the training distribution. In contrast, distribution-matching methods explicitly align the expert and policy state-action distributions, which helps mitigate distributional shift and improves robustness. Among them, Adversarial Imitation Learning (Ho and Ermon, 2016; Fu et al., 2018; Ghasemipour et al., 2020; Orsini et al., 2021)—inspired by GANs—has shown strong results but suffers from instability. Previous work has focused on stabilizing the discriminator using improved loss functions (Luo et al., 2024; Wang et al., 2024), architectures (Lai et al., 2024), and regularizers (Orsini et al., 2021; Xiao et al., 2019). However, reward assignment—how discriminator logits should be mapped to rewards—has received little attention. Fu et al. (2018) first highlighted the impact of RA, showing that performance can vary with the structure of the underlying MDP, and proposed methods such as incorporating absorbing states to mitigate this dependence. Ghasemipour et al. (2020) and Zhang et al. (2020) highlighted the impact of RA functions on policy optimization and along with Ke et al. (2021) unified various RA functions as instances of divergence minimization within the distribution-matching IL framework, while Orsini et al. (2021) empirically benchmark these functions across multiple tasks. This work is, to the best of our knowledge, the first that explores the discovery of novel reward assignment functions, as an alternative to relying on the existing human-derived ones. Recent non-adversarial distribution-matching IL methods, such as ValueDICE (Kostrikov et al., 2020) and IQ-Learn (Garg et al., 2021), have demonstrated promise but showed mixed empirical performance, with AIL variants often performing comparably or better across multiple benchmarks (Lai et al., 2024; Jain et al., 2024). Additionally, unlike AIL approaches, they provide limited flexibility in accommodating scenarios such as state-only demonstrations (Torabi et al., 2018; Jain et al., 2024) and reward shaping (Sapora et al., 2024). Another closely related area is Inverse Reinforcement Learning (Ng and Russell, 2000), which seeks to infer the underlying reward function that best explains the expert policy (Skalse and Abate, 2023; Chirra et al., 2024). This contrasts with our narrower focus on directly optimizing imitation policies and designing reward assignment functions that ensure their stable training. On the other hand, works such as (Ma et al., 2024a; b; Li et al., 2025) directly tackle the problem of reward design in MDPs by leveraging LLM-based evolution, ultimately outperforming human-designed rewards.
@@
Meta-Learning (also known as ‘learning to learn’) aims to automatically discover learning algorithms through end-to-end optimization (Schmidhuber, 1987; Thrun and Pratt, 1998) and has seen extensive application in RL (Beck et al., 2025). Historically, RL training pipelines have been constrained by CPU-bound simulators. The advent of GPU-accelerated environments, however, has enabled speedups of up to $1000\times$, revitalizing interest in Meta-RL. This has facilitated the discovery of novel loss functions (Lu et al., 2022; Jackson et al., 2024), activation functions (Nadimpalli et al., 2025), and optimizers (Goldie et al., 2024) for Deep RL. Nevertheless, differentiating through the learning process of an RL algorithm remains a challenging problem, motivating the use of black-box optimization methods (Salimans et al., 2017; Chen et al., 2023b). However, these approaches typically suffer from high sample complexity and limited interpretability. Large language models (LLMs) provide a promising alternative as their broad domain knowledge and code-generation capabilities can enable more interpretable and effective learning algorithms (Goldie et al., 2025). Beyond RL, LLM-based black box optimization has demonstrated success in a wide range of domains, including environment generation (Faldor et al., 2024), neural architecture search (Chen et al., 2023a), combinatorial optimization (Ye et al., 2024), mathematics (Romera-Paredes et al., 2024), and more (Novikov et al., 2025).
@@ 3 Background
We consider a Markov Decision Process (MDP) $\mathcal{M}$ defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,\mu)$, where $\mathcal{S}$ denotes the set of states, $\mathcal{A}$ is the set of actions, $\mathcal{P}(s^{\prime}|s,a)\in[0,1]$ is the transition probability, $r(s,a)\in\mathbb{R}$ is the reward function, $\gamma\in[0,1]$ is the discount factor, and $\mu(s)\in\Delta(\mathcal{S})$ is the initial state distribution. A policy $\pi(.|s)\in\Delta(\mathcal{A})$ is a distribution over the set of valid actions for state $s$. A trajectory $\tau=\{(s_{t},a_{t})\}$ denotes the state-action pairs encountered by executing $\pi$ in $\mathcal{M}$. For a given policy $\pi$, the occupancy measure $\rho_{\pi}(s,a)$ is defined as $\rho^{\pi}(s,a)=(1-\gamma)\mathbb{E}_{\tau\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s,a_{t}=a)$]. Intuitively, it can be interpreted as the distribution over state-action pairs that the agent encounters while following policy $\pi$. A one-to-one correspondence exists between $\pi$ and $\rho_{\pi}$ (Syed et al., 2008), allowing us to use them interchangeably. For functions $f(s,a)$ dependent only on state-action pairs, we have $\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}f(s_{t},a_{t})\right]=\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim\rho_{\pi}}[f(s,a)].$ We leverage this identity to use the two expectations interchangeably when optimizing over $f$.
@@
Jensen-Shannon GAIL (Ho and Ermon, 2016) $\text{softplus}(\ell(s,a))$
@@ 4 Problem Definition
In this paper, we focus on Step 3, highlighting that providing an informative learning signal is crucial for effective policy improvement and overall adversarial training. Originally discussed by Ghasemipour et al. (2020), Figure 1 illustrates how the reward assignment function shapes the policy’s learning dynamics. The AIRL RA function encourages the policy to visit state-action pairs where expert visitation exceeds its own and penalizes it equally when it surpasses the expert. In contrast, the GAIL RA function only incentivizes matching the expert on underrepresented pairs, while its heuristic variant does the opposite. FAIRL employs a more nuanced approach: it rewards the policy for slightly exceeding expert visitation, but imposes steep penalties when the expert dominates.  Ghasemipour et al. (2020) argues that this drives the policy to gradually expand and cover the expert distribution from the outside in. Furthermore, although existing RA functions (Table 1) are derived from well-established $f$-divergence theory, they neglect the practical stability challenges that arise during training.
This raises the question: Can we meta-learn a reward assignment function that results in stable and effective adversarial training? To evaluate training quality, we measure the divergence between the expert and the policy after training. Specifically, we use the Wasserstein distance (Rubner et al., 1998), computed between rollouts generated by the policy and expert demonstrations. We chose this metric due to its robustness and sensitivity in measuring distances between occupancy distributions in RL settings (Luo et al., 2023; Rupf et al., 2024).
@@ 5 Discovering RA functions via Evolutionary Search
Optimizing Equation 3 is challenging because it requires backpropagating gradients through the entire adversarial training loop, which is generally computationally intractable. Consequently, prior work has typically relied on black-box methods for such bilevel optimization problems (Goldie et al., 2024; Lu et al., 2022). In this work, we adopt an LLM-guided evolutionary framework that has been shown to be sample-efficient, interpretable, and capable of discovering generalizable algorithms in meta-RL (Goldie et al., 2025).
@@ 6 Empirical Studies
We begin by describing the experimental setup, after which we present the results. We conducted our experiments on two benchmark suites: MuJoCo control tasks (Ant, Reacher, Walker2d, HalfCheetah, and Hopper(Todorov et al., 2012) and Minatar (Asterix, SpaceInvaders, and Breakout(Young and Tian, 2019). For each task, we collect 10 successful expert demonstrations from a PPO-trained policy (Schulman et al., 2017) and subsample every 20th transition, following standard practice (Ho and Ermon, 2016). Unless stated otherwise, we optimize the policy using PPO and regularize the discriminator with a gradient penalty (Gulrajani et al., 2017). Consistent with prior work in IL, we adopt fixed-length episodes of 1000 timesteps for MuJoCo environments (Gleave et al., 2022). However, we do not enforce this for the Minatar tasks since doing so significantly degraded performance. Our implementation is fully written in JAX (Bradbury et al., 2018), using PureJaxRL (Lu et al., 2022), Brax (Freeman et al., 2021) for MuJoCo environments, Gymnax (Lange, 2022) for Minatar, and OTT-JAX (Cuturi et al., 2022) for computing Wasserstein distance. The complete training hyperparameters are provided in Appendix D. We normalize returns using min-max scaling between random and expert policy performance, with all results averaged over 16 independent seeds.